OpenStudy (anonymous):

3x+y =7 3x+y=10 if the system cannot be solved state that it cannot be solved and explain why... plzz help!!

OpenStudy (turingtest):

3x+y=7 3x+y=10 subtract one equation from the other, what do you get?

OpenStudy (anonymous):

3x+y=17?

OpenStudy (turingtest):

not quite subtracting one equation from the other looks like this: 3x+y=7 -(3x+y=10) ----------- (3x+y)-(3x+y)=7-10 simplify^^^^ what do you get?

OpenStudy (anonymous):

y=-3?

OpenStudy (turingtest):

closer, but the left hand side is still wrong: (3x+y)-(3x+y)=0 so it should be...?

OpenStudy (anonymous):

-3?

OpenStudy (turingtest):

you need to write the whole equation -3 is on the right, what's on the left side?

OpenStudy (anonymous):

im guessing nothing right?

OpenStudy (turingtest):

not nothing as I said (3x+y)-(3x+y)=0 and 0 is different than nothing (though it may not seem that way)

OpenStudy (anonymous):

ok so its 0=-3?

OpenStudy (turingtest):

yes :) and is that ever a true statement?

OpenStudy (anonymous):

no so this system doesnt have a solution?

OpenStudy (turingtest):

correct :) nice job!

OpenStudy (anonymous):

thank you!!

OpenStudy (turingtest):

welcome! there are faster ways to see that this has no solution perhaps jim_thompson will explain another method

jimthompson5910 (jim_thompson5910):

Another way to look at this: If we let z = 3x+y, then we go from 3x+y = 7 3x+y = 10 to z = 7 z = 10 Since z is a single number, it is impossible for it to be equal to BOTH 7 and 10 at the same time. So there are no solutions for z Consequently, this extends to the fact that there are no solutions for x and y.

OpenStudy (anonymous):

thank you!

OpenStudy (turingtest):

yes, that's the slick way to look at it :)

OpenStudy (anonymous):

lol!

jimthompson5910 (jim_thompson5910):

Another way is to use substitution Solve 3x+y = 7 for y to get y = -3x+7. Now plug this into the second equation 3x+y = 10 3x+(-3x+7) = 10 3x-3x + 7 = 10 0x + 7 = 10 7 = 10 ....which is NEVER true...so there are NO solutions Whatever method you use, you'll find that there are no solutions.